one leg of a right triangle is 12 cm longer than the other, and the area of the right triangle is 110 cm squared. Find the length of the longer leg
Answer by ewatrrr(24785)


Hi
one leg of a right triangle is 12 cm longer than the other
Let x and (x+12cm) represent the lengths of the legs of this right triangle
A = (1/2)b*h
110cm^2 = (1/2)x(x+12cm)
220 = x^2 + 12x
0 = x^2 + 12 - 220
factoring
(x+22)(x-10)= 0 Note:SUM of the inner product(22x) and the outer product(-10x) = 12x
(x+22) = 0
x = -22 extraneous solution
(x-10)= 0
x = 10 10cm and 22cm are the lengths of the two legs
Letting c represent the length of the longer lefg & applying the Pythagorean Theorem
10^2 + 22^2 = c^2
584 = c^2 (discarding negative solution for length)
24.17cm = c